OpenStudy (anonymous):

Solving 2-variable sytem by matrices -9x+13y=3 2x-3y=-1

OpenStudy (anonymous):

i have to use the matrices form :/

OpenStudy (anonymous):

Should it using matrices? but we can use the simple one, like elimination for example..

OpenStudy (anonymous):

right. but my work has to be in matrices........

OpenStudy (anonymous):

the augmented matrix is [-9 13 3] [ 2 -3 -1] you can change position

OpenStudy (anonymous):

1/-9(-3)-13(2) = [3 -13] [-2 9] 1/27-26 1/-1 = [3 -13] [-2 9] right?

OpenStudy (anonymous):

[ 2 -3 -1] new position, mult R1 by 1/2 to get [-9 13 3] [ 1 -3/2 -1/2] --> [ 1 -3/2 -1/2] [-9 13 3] 9R1+R2 -> [ 0 -1/2 -3]

OpenStudy (anonymous):

[ 2 -3 -1] =R1 row 1 or first row [-9 13 3] =R2

OpenStudy (anonymous):

mult R1 by 1/2 to get [ 1 -3/2 -1/2] mult R1 by 9 to get [ 9 -27/2 -9/2] then add thus to R2 [-9 13 3] to get -------------------- [ 0 -1/2 -3/2] so the new form now becomes [ 1 -3/2 -1/2] =R1 [ 0 -1/2 -3/2] =R2 now R2 means 0X -1Y/2 =-3/2 NOW SOLVE FOR Y FIRST Y=-2(-3/2) Y=3 SUBS Y TO R1 OR R2 OF THE FIRRST EQ

OpenStudy (anonymous):

are you getting it?

OpenStudy (anonymous):

yes. i am thank you

OpenStudy (anonymous):

are you in LA or SF or in SD?

OpenStudy (anonymous):

im in LA you?

OpenStudy (anonymous):

is this a cell phone number?

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

I wouldnt want to keep you up. lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is supposed the matrix form should be: the augmented matrix is [-9 13 3] [ 2 -3 -1] you can change position [ 2 -3 -1] R1x(1/2) --> [ 1 -3/2 -1/2] --> [ 1 -3/2 -1/2] [-9 13 3] --> [-9 13 3] 9R1+R2 --> [ 0 -1/2 -3/2] the new matrix form are [ 1 -3/2 -1/2] [ 0 -1/2 -3/2] when ever you have zero here in R2 means that 0X -1Y/2=-3/2 then you can solve for y=-2(-3/2) y=3 use this to salve for x on either -9x+13y=3 2x-3y=-1